OpenStudy (skullpatrol):

What is the most convincing argument you have found for why can you never divide by zero?

OpenStudy (anonymous):

given by my teacher xD

OpenStudy (anonymous):

I think it is because u can't divide something in nothing, u are going to do nothing, there if nothing to do!

OpenStudy (anonymous):

there is*

OpenStudy (skullpatrol):

But isn't that the same as dividing by 1? Because you have done nothing and left the number unchanged, right? x/1 = x

OpenStudy (cwrw238):

you add add as many 0's as you like - ad infinitum - but you'll always get a total of zero

OpenStudy (cwrw238):

* you can add

OpenStudy (cwrw238):

also if dividing by 0 is valid you can prove things like 1 = 2

OpenStudy (cwrw238):

if we let x = 1 then x -1 = 2(x -1) divide both sides by x-1 :- 1 = 2

OpenStudy (zzr0ck3r):

\(\mathbb{R}\) is not a group with multiplication. So the question does not even make sense. :)

OpenStudy (zzr0ck3r):

For the same reason I can't divide house by car

OpenStudy (skullpatrol):

What does "a group with multiplication" mean?

OpenStudy (zzr0ck3r):

For a set to form a group it needs to have the property that every element has an inverse. Multiplication on R does not pass this test. The identity is 1, but there is no number you can multiply by 0 to get 1. So that R with multiplication is not a group. R\{0} is however

OpenStudy (zzr0ck3r):

so in short, since there is no number we can multiply 0 by to get 1, we cant divide by 0

OpenStudy (skullpatrol):

Yes, that is convincing if you define division as multiplication by the reciprocal. But what about a/b = c iff a = b•c?

OpenStudy (zzr0ck3r):

that implies that a/b is defined, which is not for any additive identity on a ring, when we are calling the two operations additive and multiplicative.

OpenStudy (zzr0ck3r):

there are two operations on the real numbers, additive and multiplication

OpenStudy (zzr0ck3r):

division is not an operation for a group or ring because it is NOT associative

OpenStudy (zzr0ck3r):

To get a very basic understanding of why this must be, go look up groups, and rings, and then fields if you get bored.

OpenStudy (skullpatrol):

If b = 0, the above definition is still valid if a = 0, correct? But the c is no longer unique. That^ is the second most convincing argument I've seen after yours :-)

OpenStudy (zzr0ck3r):

what that really says is \(ab^{-1} = c \iff a=cb\)

OpenStudy (zzr0ck3r):

I think its best not to think of division as a thing, its really the inverse of a thing ;)

OpenStudy (skullpatrol):

Indeed, it then fits perfectly with subtraction.

OpenStudy (zzr0ck3r):

but that is not defined for 0 by properties of being a group and you need those properties to make a nice structure.

OpenStudy (amistre64):

there is no rule that says you can never divide by zero.

OpenStudy (amistre64):

some courses define 1/0 as infinity

OpenStudy (amistre64):

unless otherwise defined, dividing a real number by 0 remains 'undefined'.

OpenStudy (freckles):

how would you know if 1/0 is infinity or negative infinity ?

OpenStudy (amistre64):

when defining the slope of a line Ax + 0y = C 0y = -Ax + C y = -A/0 x + C/0 the slope of a vertical line is .... infinity for some.

OpenStudy (amistre64):

again, depends on the course ... i leave it to them to define it the way they see fit.

OpenStudy (freckles):

\[\lim_{x \rightarrow 0^+}\frac{1}{-x}\] if you plug in 0 you get 1/0 but the limit is negative infinity not positive infinity if i take that negative sign off though you would your 1/0 where the limit goes to positive infinity

OpenStudy (freckles):

oh are you defining undefined as infinity ?

OpenStudy (amistre64):

infinity is not a real number :)

OpenStudy (freckles):

oh wait no nevermind kinda confused I always seen vertical lines with undefined slopes

OpenStudy (freckles):

i know infinity isn't a number but just because something isn't defined doesn't mean it is infinity

OpenStudy (amistre64):

1/0 is not a real number, and x+1 = x has no real solution

OpenStudy (freckles):

or even negative infinity it could just not be defined at all

OpenStudy (perl):

Some books say we can't define a/0 because if a/0 = k, then a = 0*k and this leads to a contradiction. for example if 3/0 = k, then 3 = 0*k , but then 3 = 0. this is supported by the first post here http://mathforum.org/dr.math/faq/faq.divideby0.html But the problem with this line of reasoning is, 0/0= 0 seems to be valid, since 0 = 0*0. so you really need a more sophisticated answer than that.

OpenStudy (amistre64):

if we define it, then it is defined. I havent taken the courses so I wouldnt be all that knowledgable about the specfic.

OpenStudy (amistre64):

dividing by zero is undefined, until it is defined. if it comports itself to a noncontradictory philosophy, then its valid.

OpenStudy (freckles):

oh so I guess you are saying it is up to the instructor to define a/0 in my opinion a/0 is indeterminate you cannot determine it as infinity or -negative infinity I prefer to say a/0 is undefined because of that until we get to limits of course when we actually consider what something is approaching or whatever but even then sides might not agree and you can't say anything but it is undefined like \[\lim_{x \rightarrow 0}\frac{1}{-x} \text{ is undefined since both sides don't agree on the same infinity }\]

OpenStudy (freckles):

so it just seems kinda confusing to define 1/0 as infinity when students get to calculus they will see different stories to the 1/0 thing

OpenStudy (freckles):

but anyways I'm done about division by 0 this topic is scary

OpenStudy (perl):

in calculus we say expressions 0/0 (in limits) are indeterminate, but a/0 with a not equal to zero diverge

OpenStudy (perl):

indeterminate has a special meaning in the context of limits, it is not the same as undefined

OpenStudy (freckles):

I'm sorry what i meant by what I said a/0 can't be determined as infinity or negative infinity should have just said that even though that still seems not determined :p

OpenStudy (perl):

no need to apologize :)

OpenStudy (perl):

i agree with that , there might be some special number system that has only one infinity

OpenStudy (freckles):

I apologize all the time when I don't mean it it is hard to tell when I really mean it and when I don't My husband or wife (that is indeterminate also) says I say sorry too much and I tell him I don't mean it half the time :p

OpenStudy (perl):

If you are pressed to, you can define a/0 with a ≠ 0 as |∞|

OpenStudy (freckles):

I think I would be more comfortable defining |a/0| where a not equal 0 as infinity

OpenStudy (skullpatrol):

But using absolute value bars around infinity is treating it like a real number?

OpenStudy (perl):

yes that makes sense

OpenStudy (perl):

you can treat infinity as a real number in something called the extended Real life

OpenStudy (skullpatrol):

life?

OpenStudy (freckles):

lol line

OpenStudy (perl):

line*

OpenStudy (skullpatrol):

:D great discussion, thanks.

OpenStudy (kainui):

I disagree. Dividing my zero is just something you have to handle carefully and that's what calculus is. Even in projective geometry we can safely define infinity as a point that you approach from all directions in the plane for instance and it makes perfect geometric sense that -infinity and +infinity are the same thing! I'll draw a picture if anyone's curious.

OpenStudy (freckles):

I'm curious. When I think about the opposite infinities, I think about two totally different directions.

OpenStudy (freckles):

Are you saying it is kinda circular or something?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @zzr0ck3r so in short, since there is no number we can multiply 0 by to get 1, we cant divide by 0 \(\color{blue}{\text{End of Quote}}\) \[\huge\rm (0)^0 = 1 \] lel :D ;P

OpenStudy (anonymous):

all X^0=1 except for 0^0=0

OpenStudy (kainui):

Yeah, so I'll start with the 2D version and the 3D or higher version is not too crazy. In projective geometry we're just saying we're taking every point on the line and projecting it to a point on the circle:  When the projected line is parallel to the top of the circle, that point is called infinity and if we tilt it either way it will either give us a negative or positive number. The plane version is essentially the same thing, but still only a singular infinite point at the top of the sphere.  so if you pick any point infinitely far away it will always point to the exact same point at the top, and it's nice to have for doing computer graphics apparently to have a point at infinity or so I hear, I was more interested in it for other reasons so I didn't get into any applications of it.